OpenStudy (anonymous):

integral x/((1-x^2)^(1/2)) dx

OpenStudy (anonymous):

\[\int\limits \frac{x}{\sqrt{1-x^2}}dx ?\]

OpenStudy (anonymous):

Make the substitution: \[\zeta=1-x^2 \implies d \zeta=-2x dx \implies \frac{-1}{2} d \zeta =x dx\] Replacing this the integral becomes: \[\frac{-1}{2}\int\limits \frac{d \zeta}{\sqrt{\zeta}}=\frac{-1}{2}\int\limits \zeta^{\frac{-1}{2}}d \zeta=-\zeta^{\frac{1}{2}}+C=-\sqrt{1-x^2}+C\]

OpenStudy (anonymous):

Here is the answer, does it make sense?